# A Szegö limit theorem for a class of Toeplitz operators on the Bergman space of the unit ball with singular symbols Daniel Ivan Ramirez Montaño

### v1

@@
Daniel Ivan Ramirez Montaño
We obtain a Szegö limit theorem for a family of Toeplitz operators defined on the weighted Bergman space of the unit ball $\mathbb{B}_{n}$. The symbols of these operators are supported on some isotropic or co-isotropic submanifold $\Gamma\subseteq\mathbb{B}_{n}$ and can be seen, in general, as measures that are singular with respect to the Lebesgue measure on $\mathbb{C}^{n}$.
@@ 1 Introduction
where $\lambda_{j,N}$ are the eigenvalues of the $N$-th order truncation of the Toeplitz matrix whose diagonals are the Fourier coefficients of a real valued $a\in L^{\infty}(\mathbb{S}^{1})$ and $f$ is a continuous function on $[\mathrm{ess}\inf a,\mathrm{ess}\sup a]$. Later works focused on reproducing this result with weaker conditions for $a$ and applying to different families of functions $f$ (e.g. [14], [16]).
Some of these ideas were generalized in [3]. For a bounded self-adjoint operator $A$ on a Hilbert space $H$, we want a measure $\mu_{A}$ on $\mathbb{R}$ such that
@@
for any $f\in C(\mathbb{R})$. If $H$ has finite dimension, this can be done using the Riesz-Markov-Kakutani theorem, since the assignation $f\mapsto\mathrm{Tr}(f(A))$ results in a continuous functional on $C(\mathbb{R})$. In fact, $\mu_{A}$ is the sum of the Dirac measures centered on each eigenvalue of $A$. In general, this cannot be done when $H$ is not of finite dimension, so an alternative construction is proposed. Taking an increasing sequence of finite-rank orthogonal projections $P_{n}$ on $H$ such that $P_{n}$ converges to the identity in $H$, the correspondent sequence of measures $\mu_{n}$ satisfying (1) for the operator $P_{n}A|_{P_{n}(H)}$ is constructed. Therefore, a natural substitute for $\mu_{A}$ would be the weak limit of the $\mu_{n}$ measures. In this situation a Szegö limit theorem translates to finding a measure whose integral reproduces the limit of the traces of continuous images of a sequence of operators. Particular attention has been paid to this problem when $A$ is replaced by a pseudo-differential operator (see [3], [8], [15]).
Recently in [2] results of this type were obtained considering various function spaces. In particular, the weighted Bergman space of the disc $\mathbb{D}$ and the weighted Fock space on $\mathbb{C}$ are treated. In [11] the problem was studied in the context of the Fock space of $\mathbb{C}^{n}$ and considering symbols supported on some manifold, thus in general being measures that are singular with respect to the Lebesgue measure. In both of these works, the family of operators whose limiting behavior is considered is indexed by the parameter defining the associated weighted space. We will also follow this approach, obtaining an analogue result to the main one of [11] for the case of the weighted Bergman space of the unit ball $\mathbb{B}_{n}$.
@@ 1.1 Definitions
For $n\in\mathbb{N}$, let $\mathbb{B}_{n}:=\{z\in\mathbb{C}^{n}:|z|<1\}$ denote the unit ball in $\mathbb{C}^{n}$. For $\alpha>-1$, the Bergman space with standard weights $A^{2}_{\alpha}(\mathbb{B}_{n})$ is the Hilbert space of all holomorphic functions in $L^{2}(\mathbb{B}_{n},\mathrm{d}v_{\alpha})$, where
@@
When $\alpha=0$, $\mathrm{d}v_{0}=\mathrm{d}v$ and $A^{2}(\mathbb{B}_{n})$ and $K(z,w)$ are used for $A^{2}_{0}(\mathbb{B}_{n})$ and $K^{0}(z,w)$, respectively. These are the standard (unweighted) Bergman spaces. The function $K^{\alpha}$ is known as the (weighted) Bergman kernel and it satisfies the reproducing property
$f(z)$ $=$ $\big{\langle}f,K^{\alpha}(\ \cdot\ ,z)\big{\rangle}_{A^{2}_{\alpha}(\mathbb{B}_{n})}$
$=$ $\int_{\mathbb{B}_{n}}K^{\alpha}(z,w)f(w)\ \mathrm{d}v_{\alpha}(w)\ ,$
for any $f\in A^{2}_{\alpha}(\mathbb{B}_{n})$, $z\in\mathbb{B}_{n}$. This last integral expression makes sense even for more general classes of functions. In fact, it represents the orthogonal projection from $L^{2}(\mathbb{B}_{n},\mathrm{d}v_{\alpha})$ to $A^{2}(\mathbb{B}_{n})$, also called the Bergman projection. For a bounded Lebesgue measurable funtion $a$ or a regular Borel measure $\mu$, both on $\mathbb{B}_{n}$, their Toeplitz operators $\mathcal{T}_{a}$ and $\mathcal{T}_{\mu}$ are defined by
@@
for any $f\in A^{2}(\mathbb{B}_{n})$, $z\in\mathbb{B}_{n}$. In each case, $a$ or $\mu$ is called the symbol of the respective Toeplitz operator. The operator $\mathcal{T}_{\mu}$ is a natural generalization of $\mathcal{T}_{a}$, which in turn is also defined as being the compression to $A^{2}(\mathbb{B}_{n})$ of the multiplying operator $M_{a}f=af$.
We present the Toeplitz operators we will work with.
@@
Let $\Gamma\subset\mathbb{B}_{n}$ be a Riemannian submanifold of dimension $d$, $\mathrm{d}\sigma$ the Riemannian volume element on $\Gamma$, and $a:\Gamma\rightarrow\mathbb{C}$ a function such that $a\in C^{2}_{0}(\Gamma)$. We define the Toeplitz operator $T_{a\mathrm{d}\sigma}$ by
@@
  • Although not made explicit in the notation, $T_{a\mathrm{d}\sigma}$ depends on $\alpha$. We are interested in the behavior of $T_{a\mathrm{d}\sigma}$ as $\alpha$ tends to infinity.
  • If $d<2n$, the resulting symbol measure $\mu=a\mathrm{d}\sigma$ is singular with respect to the Lebesgue measure on $\mathbb{B}_{n}$.
@@
  • As $a$ is taken to have compact support, $T_{a\mathrm{d}\sigma}$ is bounded, compact, and belongs to each Schatten class $S_{p}(A_{\alpha}^{2}(\mathbb{B}_{n}))$, $p>0$.
  • The condition $a\in C^{2}(\Gamma)$ can be changed to $a\in L^{\infty}(\Gamma)$ without affecting the definition of $T_{a\mathrm{d}\sigma}$. In fact, it is not needed for a substantial part of the results. However, it is included in the definition since it is necessary for the main theorem.
@@
By writing the local holomorphic coordinates as $z_{\ell}=x_{\ell}+iy_{\ell}$, $\ell=1,{\scriptstyle\dots},n$, $\mathbb{B}_{n}$ can be naturally regarded both as a complex manifold of dimension $n$ and as a smooth manifold of dimension $2n$. For any $p\in\mathbb{B}_{n}$ the respective real and complex tangent spaces are given in terms of basic tangent vectors by
@@
with the relation between the elements of these bases being given by the Wirtinger derivatives
@@
for $\xi,\nu\in\mathrm{T}_{p}\mathbb{B}_{n}$.
Let $V\subseteq\mathrm{T}_{p}\mathbb{B}_{n}$ be a subspace and denote by
@@
We say that $V$ is isotropic (respectively co-isotropic, respectively Lagrangian) if $V\subseteq V^{\omega}$ (respectively $V\supseteq V^{\omega}$, respectively $V=V^{\omega}$).
@@
From the skew-symmetry of $\omega_{p}$, it is straightforward that any subspace $V$ of dimension one is isotropic. On the other hand, if $V$ has codimension one then it is co-isotropic, as in this case $V^{\omega}$ is isotropic and so $V^{\omega}\subseteq(V^{\omega})^{\omega}=V$. In addition, the total subspace $\mathrm{T}_{p}\mathbb{B}_{n}$ is trivially co-isotropic.
@@
It follows from Example 1.5 that any curve is isotropic and any hypersurface is co-isotropic. Also, $\mathbb{B}_{n}$ or any open subset of $\mathbb{B}_{n}$ is co-isotropic.
@@
$d^{\prime}=\left\{\begin{array}[]{cl}d&\Gamma\mbox{ \ isotropic}\\ 2n-d&\Gamma\mbox{ \ co-isotropic}\end{array}\right.\ .$
@@ 1.2 Statement of the main result
[Szegö limit theorem] Let $\Gamma\subseteq\mathbb{B}_{n}$ be a isotropic or co-isotropic submanifold, $a\in C^{2}_{0}(\Gamma)$ a positive function, and $R>0$ such that $\sup\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}<R$. If $\varphi$ is a function such that $\frac{\varphi(t)}{t^{p}}\in C[0,R]$ for some $p>0$, then
@@
We single out the following particular situation. A form of case $n=1$ appears as Theorem 1.1 in [2].
@@
The case $\Gamma=\mathbb{B}_{n}$ is trivially co-isotropic and $d^{\prime}=0$. We thus get
@@
The article has the following structure. In the following section, we describe some aspects of the intrinsic geometry of $\Gamma$. We obtain a relation involving the Riemannian metric and the symplectic form in $\Gamma$ that will be needed later. In Section 3 we derive an integral representation for the trace of a composition of multiple Toeplitz operators with symbols sharing the same support, and then in Section 4 we make use of the stationary phase method to study the asymptotic behavior of this integral as $\alpha$ grows to infinity. The expression we obtain relies on computing the determinant of a given Hessian matrix, which we leave for Appendix B. The proof for the Szegö limit theorem is delivered in Section 5, first for polynomials and then for a wider family of continuous functions. A key step in the proof involves bounding the traces for small powers (exponents $10<p<$ 1) of the Toeplitz operators. This is done in Appendix A. Section 6 is reserved for establishing some corollaries to the main theorem, involving results on eigenvalue distribution and Schatten norm estimations.
The line of reasoning used here follows that in [11], with some of the proofs presented being adaptations to the Bergman setting of the ones therein. In particular, the proofs in Section 6 and Appendix B.1 remain virtually unmodified but are regardless included for the sake of completeness. On the other hand, parts of appendices A and B.2 generalize their analogues there.
Throughout the article, we will use the lone capital letter $C$ to represent an arbitrary constant that may change from line to line. When deemed appropriate, we will highlight its independence from other variables.
@@ 2 The Riemannian metric on $\Gamma$
Let $\gamma:B(0,1)\rightarrow\Gamma$ be a parametrization. For each ${p=\gamma(t)\in\Gamma}$, $t=(t_{1},{\scriptstyle\dots},t_{d})\in B(0,1)$, the tangent space $\mathrm{T}_{p}\Gamma$ is spanned by the tangent vectors $\frac{\partial}{\partial\gamma_{j}}\Big{\rvert}_{t}$, $j=1,{\scriptstyle\dots},d$. These tangent vectors are defined by
$\frac{\partial}{\partial\gamma_{j}}\Big{\rvert}_{t}:$ $C^{\infty}(\Gamma)$ $\rightarrow\mathbb{R}$
$f$ $\mapsto\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}f\big{(}\gamma(t_{1},{\scriptstyle\dots},t_{j}+s,{\scriptstyle\dots},t_{d})\big{)}\Big{]}\Bigr{\rvert}_{s=0}\ .$
Also recall that for each $v\in\mathrm{T}_{p}\Gamma$, $\mathrm{D}_{p}\iota\left(v\right)$ is the tangent vector to $\mathbb{B}_{n}$ at $p=\iota(p)$ determined by
$\mathrm{D}_{p}\iota\left(v\right)(f)=v(f\circ\iota)\ ,\ \ f\in C^{\infty}\left(\mathbb{B}_{n}\right)\ .$
As a submanifold $\Gamma\subseteq\mathbb{B}_{n}$, the Riemannian metric $g$ in $\Gamma$ is given by the pullback of the metric $\mathring{b}$ in $\mathbb{B}_{n}$. More precisely
@@
$\forall\ u,v\ \in\mathrm{T}_{p}\Gamma$. We also define a skew-symmetric, nondegenerate bilinear form by
$h_{p}(u,v)=\omega_{p}\big{(}\mathrm{D}_{p}\iota(u),\left(\mathrm{D}_{p}\iota(v)\right)\big{)}\ ,$
@@
The following lemma relating $g_{p}$ and $h_{p}$ looks rather technical, but will be key during Section 4.
@@
from which (9) is also deduced.

### v2

@@
Daniel Ivan Ramirez Montaño [^1] Instituto de Matemáticas, Unidad Cuernavaca, Universidad Nacional Autónoma de México. E-mail address: daniel.ramirez@im.unam.mx
We obtain a Szegö limit theorem for a family of Toeplitz operators defined on the weighted Bergman space of the unit ball $\mathbb{B}_{n}$. The symbols of these operators are supported on some isotropic or co-isotropic submanifold $\Gamma\subseteq\mathbb{B}_{n}$ and can be seen, in general, as measures that are singular with respect to the Lebesgue measure on $\mathbb{C}^{n}$. The given theorem allows to describe the asymptotic behavior of these operators as the parameter of the weighted Bergman space tends to infinity.
@@ 1 Introduction
where $\lambda_{j,N}$ are the eigenvalues of the $N$-th order truncation of the Toeplitz matrix whose diagonals are the Fourier coefficients of a real valued $a\in L^{\infty}(\mathbb{S}^{1})$ and $f$ is any continuous function on $[\mathrm{ess}\inf a,\mathrm{ess}\sup a]$. Later works focused on reproducing this result with weaker conditions for $a$ and applying to different families of functions $f$ (e.g. [14], [16]).
The generalization of some of these ideas to the operator setting is described in [3]. For a bounded self-adjoint operator $A$ on a Hilbert space $H$, we want a measure $\mu_{A}$ on $\mathbb{R}$ such that
@@
for any $f\in C(\mathbb{R})$. If $H$ has finite dimension, this can be done using the Riesz-Markov-Kakutani representation theorem, since the assignation $f\mapsto\mathrm{Tr}(f(A))$ results in a continuous functional on $C(\mathbb{R})$. In fact, $\mu_{A}$ is the sum of the Dirac measures centered on each eigenvalue of $A$. In general, this cannot be done when $H$ is not of finite dimension, so an alternative construction is proposed. Taking an increasing sequence of finite-rank orthogonal projections $P_{n}$ on $H$ such that $P_{n}$ converges to the identity in $H$, the correspondent sequence of measures $\mu_{n}$ satisfying (1) for the operator $P_{n}A|_{P_{n}(H)}$ is constructed. Therefore, a natural substitute for $\mu_{A}$ would be the weak limit of the $\mu_{n}$ measures. In this situation a Szegö limit theorem translates to finding a measure whose integral reproduces the limit of the traces of continuous images of a sequence of operators. Particular attention has been paid to this problem when $A$ is replaced by a pseudo-differential operator (see [3], [8], [15]).
An interesting scenario occurs when considering a weighted Hilbert space of complex functions as the dominion of the Toeplitz operators. By varying the parameter that defines the associated weights we obtain a parametrized family of operators whose limiting behavior can be studied. This idea was followed in both [2] and [11]. In [2] results of this type were obtained considering Toeplitz operators on the Bergman space of the disc $\mathbb{D}$ and the Fock space of $\mathbb{C}$. Previously in [11] the problem was studied in the context of the Fock space of $\mathbb{C}^{n}$ and considering symbols supported on some submanifold of $\mathbb{C}^{n}$, thus in general being measures that are singular with respect to the Lebesgue measure. We will also follow this approach, obtaining an analogue result to the main one of [11] for the case of the weighted Bergman space of the unit ball $\mathbb{B}_{n}$.
@@ 1.1 Definitions
For $n\in\mathbb{N}$, let $\mathbb{B}_{n}:=\{z\in\mathbb{C}^{n}:|z|<1\}$ denote the unit ball in $\mathbb{C}^{n}$. For $\alpha>-1$, the weighted Bergman space $A^{2}_{\alpha}(\mathbb{B}_{n})$ is the Hilbert space of all holomorphic functions in $L^{2}(\mathbb{B}_{n},\mathrm{d}v_{\alpha})$, where
@@
When $\alpha=0$, $\mathrm{d}v_{0}=\mathrm{d}v$ and we write $A^{2}(\mathbb{B}_{n})$ and $K(z,w)$ instead of $A^{2}_{0}(\mathbb{B}_{n})$ and $K^{0}(z,w)$, respectively. The function $K^{\alpha}$ is known as the (weighted) Bergman kernel and it satisfies the reproducing property
$f(z)$ $=$ $\int_{\mathbb{B}_{n}}K^{\alpha}(z,w)f(w)\ \mathrm{d}v_{\alpha}(w)\ ,$
for any $f\in A^{2}_{\alpha}(\mathbb{B}_{n})$, $z\in\mathbb{B}_{n}$. This integral expression is used to define Toeplitz operators. For a bounded Lebesgue measurable funtion $a$ or a regular Borel measure $\mu$, both on $\mathbb{B}_{n}$, their Toeplitz operators $\mathcal{T}_{a}$ and $\mathcal{T}_{\mu}$ are defined by
@@
for any $f\in A^{2}(\mathbb{B}_{n})$, $z\in\mathbb{B}_{n}$. In each case, $a$ or $\mu$ is called the symbol of the respective Toeplitz operator. Definition (2) is the classical for a Toeplitz operator and is equivalent to the compression to $A^{2}_{\alpha}(\mathbb{B}_{n})$ of the multiplying operator $M_{a}f=af$, while $\mathcal{T}_{\mu}$ is a natural generalization of $\mathcal{T}_{a}$, as $\mathcal{T}_{a}$ is the special case when $\mu$ is absolutely continuous with respect to the Lebesgue measure and $\mathrm{d}\mu=a\mathrm{d}v$.
We present the class of Toeplitz operators we will work with.
@@
Let $\Gamma\subset\mathbb{B}_{n}$ be a Riemannian submanifold of dimension $d$, $\mathrm{d}\sigma$ the Riemannian volume element on $\Gamma$, and $a:\Gamma\rightarrow\mathbb{C}$ a function such that $a\in C^{\infty}_{0}(\Gamma)$. We define the Toeplitz operator $T_{a\mathrm{d}\sigma}$ by
@@
  • Although not made explicit in the notation, $T_{a\mathrm{d}\sigma}$ depends on $\alpha$. We are interested in the asymptotic behavior of $T_{a\mathrm{d}\sigma}$ as $\alpha$ tends to infinity.
@@
  • If $d<2n$, the resulting symbol measure $\mu=a\mathrm{d}\sigma$ is singular with respect to the Lebesgue measure on $\mathbb{B}_{n}$. Therefore, $T_{a\mathrm{d}\sigma}$ is of the form (3) and can not be written as (2).
  • As $a$ has compact support, $T_{a\mathrm{d}\sigma}$ is bounded, compact, and belongs to each Schatten class $S_{p}(A_{\alpha}^{2}(\mathbb{B}_{n}))$, $p>0$.
  • Condition $a\in C^{\infty}(\Gamma)$ can be changed to $a\in L^{\infty}(\Gamma)$ without affecting the definition of $T_{a\mathrm{d}\sigma}$. In fact, it is not needed for a substantial part of the results. However, it is included in the definition since it is necessary for the main theorem.
@@
By writing the local holomorphic coordinates as $z_{\ell}=x_{\ell}+iy_{\ell}$, $\ell=1,{\scriptstyle\dots},n$, $\mathbb{B}_{n}$ can be naturally regarded both as a complex manifold of dimension $n$ and as a smooth manifold of dimension $2n$. For $p\in\mathbb{B}_{n}$ the respective real and complex tangent spaces are given in terms of basic tangent vectors by
@@
with the relation between the elements of these bases given by the Wirtinger derivatives
@@
for $\xi,\nu\in\mathrm{T}_{p}\mathbb{B}_{n}$. If $V\subseteq\mathrm{T}_{p}\mathbb{B}_{n}$ is a subspace, we denote by
@@
We say that $V$ is isotropic (respectively co-isotropic, Lagrangian) if $V\subseteq V^{\omega}$ (respectively $V\supseteq V^{\omega}$, $V=V^{\omega}$).
@@
From the skew-symmetry of $\omega_{p}$ it follows that any subspace of dimension one is isotropic and any of co-dimension one is co-isotropic. In addition, the total subspace $\mathrm{T}_{p}\mathbb{B}_{n}$ is trivially co-isotropic.
@@
Following Example 1.5, any curve is isotropic and any hypersurface is co-isotropic. Also, $\mathbb{B}_{n}$ or any open subset of $\mathbb{B}_{n}$ is co-isotropic.
1.2 Statement of the main result
@@
$d^{\prime}:=\left\{\begin{array}[]{cl}d&\Gamma\mbox{ \ isotropic}\\ 2n-d&\Gamma\mbox{ \ co-isotropic}\end{array}\right.\ .$
@@
[Szegö limit theorem] Let $\Gamma\subseteq\mathbb{B}_{n}$ be a isotropic or co-isotropic submanifold, $a\in C^{\infty}_{0}(\Gamma)$ a positive function, and $R>0$ such that $\sup\limits_{\xi\in\Gamma}\frac{a(\xi)}{(1-|\xi|^{2})^{n+1}}<R$. If $\varphi$ is a function such that $\frac{\varphi(t)}{t^{p}}\in C[0,R]$ for some $p>0$, then
@@
The case $\Gamma=\mathbb{B}_{n}$ is trivially co-isotropic, with (7) becoming
@@
The case $n=1$ appeared recently as Theorem 1.1 in [2] without the smoothness hypothesis for the symbol $a$, but for a narrower family of functions $\varphi$.
The article has the following structure. In the following section, we describe some aspects of the intrinsic geometry of $\Gamma$. We obtain a relation involving the Riemannian metric and the symplectic form in $\Gamma$ that will be needed later. In Section 3 we derive an integral representation for the trace of a composition of multiple Toeplitz operators with symbols sharing the same support and then apply the method of stationary phase to describe the asymptotic behavior of this integral as $\alpha$ grows to infinity. The expression we obtain relies on computing the determinant of a given Hessian matrix, which we leave for Appendix B. The proof of the Szegö limit theorem is delivered in Section 4, first for polynomials and then for a wider family of continuous functions. A key step in the proof involves bounding the traces for small powers (exponents $10<p<$ 1) of the Toeplitz operators. This is done in Appendix A. Section 5 is reserved for presenting some corollaries to the main theorem regarding eigenvalue distribution and Schatten norm estimates. We also expose an example illustrating the given results.
The line of reasoning used here follows that in [11], with some of the proofs presented being adaptations to the Bergman setting of the ones therein. In particular, some proofs in Section 5 and Appendix B.1 proceed virtually unmodified and are therefore omitted for the sake of briefness.
Throughout the article, we will use the lone capital letter $C$ to represent an arbitrary constant that may change from line to line. When deemed appropriate, we will highlight its independence from other variables. The symbol $B(0,\varepsilon)$ will denote the euclidean ball in $\mathbb{R}^{d}$ with radius $\varepsilon>0$ centered at the origin.
@@ 2 The Riemannian metric on $\Gamma$
Recall that for each $v\in\mathrm{T}_{p}\Gamma$, the differential of the inclusion $\iota:\Gamma\rightarrow\mathbb{B}_{n}$ induces a tangent vector $\mathrm{D}_{p}\iota(v)$ to $\mathbb{B}_{n}$ at $p=\iota(p)$. As a submanifold $\Gamma\subseteq\mathbb{B}_{n}$, the Riemannian metric $g$ in $\Gamma$ is given by the pullback of the metric $\mathring{b}$ in $\mathbb{B}_{n}$. More precisely
@@
$\forall\ u,v\ \in\mathrm{T}_{p}\Gamma$. We also define a skew-symmetric, nondegenerate bilinear form $h$ in terms of the symplectic form $\omega_{p}$ by
$h_{p}(u,v)=\omega_{p}\big{(}\mathrm{D}_{p}\iota(u),\mathrm{D}_{p}\iota(v)\big{)}\ ,$
@@
Let $\gamma:B(0,1)\rightarrow\Gamma$ be a parametrization, so that for each ${p=\gamma(t)\in\Gamma}$, the tangent space $\mathrm{T}_{p}\Gamma$ is spanned by the tangent vectors $\frac{\partial}{\partial\gamma_{j}}\Big{\rvert}_{t}$, $j=1,{\scriptstyle\dots},d$. Denote $\gamma_{\ell}:B(0,1)\rightarrow\mathbb{C}$, $\ell=1,\dots,n$ its component functions. The following lemma relating $g_{p}$ and $h_{p}$ looks rather technical, but will be key in Section 3.2.
@@
from where (9) is also deduced.